# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ VII-A Auxiliary lemmas
$\textstyle\leq 4m\sigma_{f}^{2}(\lambda_{z}^{t})^{2}t^{-1}+4\bar{L}_{f}^{2}(\lambda_{z}^{t})^{2}\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t-1}\|^{2}]$
@@
Following an argument similar to the derivation of (37), we incorporate the preceding inequality and (49) into the second term on the right hand side of (45) to establish Lemma 7. ∎
@@
Under Assumptions 1-4, if the decaying rates of stepsizes satisfy $01>v_{x}>v_{z}>v_{y}>$ 0, then we have
$\textstyle\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]$ $\textstyle\leq\frac{C_{x}}{(t+1)^{2\varsigma_{x}}},$ (50)
$\textstyle\mathbb{E}[\|\boldsymbol{y}^{t+1}-\boldsymbol{y}^{t}-\boldsymbol{1}_{m}\otimes(\bar{y}^{t+1}-\bar{y}^{t})\|^{2}]$ $\textstyle\leq\frac{C_{y}}{(t+1)^{2\varsigma_{y}}},$ (51)
$\textstyle\mathbb{E}[\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}-\boldsymbol{1}_{m}\otimes(\bar{z}^{t+1}-\bar{z}^{t})\|^{2}]$ $\textstyle\leq\frac{C_{z}}{(t+1)^{\beta_{z}}},$ (52)
where $\beta_{z}$ is given by $\beta_{z}=\min\{2v_{z}+2\varsigma_{y}-2v_{y},2\varsigma_{z}\}$ and $C_{x}$, $C_{y}$, and $C_{z}$ are given in (56), (57), and (58), respectively.
@@
$\textstyle\sum_{i=1}^{m}\!\!\left(\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t+1}\|^{2}]\!+\!\frac{|\rho_{2}|}{2}\!\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]\right)$
@@
$\textstyle\leq\left(1-\frac{|\rho_{2}|}{2}\right)\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]$
$\textstyle+(c_{y1}(\lambda_{y}^{t})^{2}+c_{z1}(\lambda_{z}^{t})^{2})\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t-1}\|^{2}]$
@@
$\textstyle\sum_{i=1}^{m}\!\!\left(\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t+1}\|^{2}]\!+\!\frac{|\rho_{2}|}{2}\!\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]\right)$
@@
$\textstyle\leq\left(1-\frac{|\rho_{2}|}{2}\right)\Big{[}\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]$
$\textstyle\quad+\frac{|\rho_{2}|}{2}\!\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]$
@@
$\textstyle\sum_{i=1}^{m}\!\!\left(\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t+1}\|^{2}]\!+\!\frac{|\rho_{2}|}{2}\!\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]\right)$ (54)
@@
where $\beta_{0}=\min\{2v_{z}+2\varsigma_{y}-2v_{y},2\varsigma_{x},2\varsigma_{y},2\varsigma_{z}\}$ and $c_{1}\!=\!b_{0}(\frac{4\beta_{0}}{e\ln(\frac{4}{4-|\rho_{2}|})})^{\beta_{0}}(\frac{a_{0}(2-|\rho_{2}|)}{2b_{0}}\!+\!\frac{4}{|\rho_{2}|})$ with $a_{0}\triangleq\sum_{i=1}^{m}\!(\mathbb{E}[\|\boldsymbol{x}_{(i)}^{1}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{1}\|^{2}]+\frac{|\rho_{2}|}{2}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{0}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{0}\|^{2}])+(1-\frac{|\rho_{2}|}{4})(\mathbb{E}[\|\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0}\|^{2}]+\mathbb{E}[\|\boldsymbol{\hat{z}}^{1}-\boldsymbol{\hat{z}}^{0}\|^{2}])$ and $b_{0}=c_{x2}+c_{y3}+c_{z4}+2mr$.
@@
$\textstyle\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t+1}\|^{2}]$ (55)
$\textstyle\leq(1-|\rho_{2}|)\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]$
@@
$\vspace{-0.2em}\textstyle\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]\leq\frac{C_{x}}{(t+1)^{2\varsigma_{x}}},\vspace{-0.2em}$ (56)
where $C_{x}$ is given by $C_{x}\!=\!b_{1}(\frac{8\varsigma_{x}}{e\ln(\frac{2}{2-|\rho_{2}|})})^{2\varsigma_{x}}(\frac{a_{1}(1-|\rho_{2}|)}{b_{1}}+\frac{2}{|\rho_{2}|})$ with $a_{1}\!=\!\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{0}-\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{0}\|^{2}]$, $b_{1}\!=\!\frac{4c_{1}c_{x1}(\lambda_{x}^{0})^{2}}{(4-|\rho_{2}|)(\lambda_{z}^{0})^{2}}+c_{x2}\sigma_{x}^{2}$, $c_{1}$ given in (54), and $c_{x1}$ and $c_{x2}$ given in (23).
@@
$\vspace{-0.2em}\textstyle\mathbb{E}[\|\boldsymbol{\hat{y}}^{t+1}-\boldsymbol{\hat{y}}^{t}\|^{2}]\leq\frac{C_{y}}{(t+1)^{2\varsigma_{y}}},\vspace{-0.2em}$ (57)
with $C_{y}=b_{2}(\frac{8\varsigma_{y}}{e\ln(\frac{2}{2-|\rho_{2}|})})^{2\varsigma_{y}}(\frac{a_{2}(1-|\rho_{2}|)}{b_{2}}+\frac{2}{|\rho_{2}|})$, $a_{2}\!=\!\textstyle\mathbb{E}[\|\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0}\|^{2}]$, $b_{2}\!=\!C_{x}c_{y1}(\lambda_{y}^{0})^{2}+c_{y3}\sigma_{y}^{2}+\frac{4c_{1}c_{y2}(\lambda_{y}^{0})^{2}(\lambda_{x}^{0})^{2}}{(4-|\rho_{2}|)(\lambda_{z}^{0})^{2}}$, $c_{1}$ given in (54), $C_{x}$ given in (56), and $c_{y1}$ through $c_{y3}$ given in (31).
@@
$\vspace{-0.2em}\textstyle\mathbb{E}[\|\boldsymbol{\hat{z}}^{t+1}-\boldsymbol{\hat{z}}^{t}\|^{2}]\leq\frac{C_{z}}{(t+1)^{\min\{2v_{z}+2\varsigma_{y}-2v_{y},2\varsigma_{z}\}}}=\frac{C_{z}}{(t+1)^{\beta_{z}}},\vspace{-0.2em}$ (58)
with $C_{z}=b_{3}(\frac{4\beta_{z}}{e\ln(\frac{2}{2-|\rho_{2}|})})^{4\beta_{z}}(\frac{a_{3}(1-|\rho_{2}|)}{b_{3}}+\frac{2}{|\rho_{2}|})$, $a_{3}\!=\!\textstyle\mathbb{E}[\|\boldsymbol{\hat{z}}^{1}-\boldsymbol{\hat{z}}^{0}\|^{2}]$, $b_{3}\!=\!C_{x}c_{z1}(\lambda_{z}^{0})^{2}\!+\!\frac{2C_{y}c_{z2}(\lambda_{z}^{0})^{2}}{(\lambda_{y}^{0})^{2}}\!+\!\frac{4c_{1}c_{z3}(\lambda_{x}^{0})^{2}}{(4-|\rho_{2}|)}+\frac{c_{z4}(\lambda_{z}^{0})^{2}\sigma_{y}^{2}}{(\lambda_{y}^{0})^{2}}+2mr\sigma_{z}^{2}$, $c_{1}$ given in (54), $C_{x}$ given in (56), $C_{y}$ given in (57), and $c_{z1}$ through $c_{z3}$ given in (44). ∎
@@
$\textstyle u_{i}^{t}=\nabla_{x}f_{i}^{t}(x_{i}^{t},\tilde{y}_{i}^{t})-\nabla g_{i}^{t}(x_{i}^{t})\tilde{z}_{i}^{t};$ (59)
$\textstyle\breve{u}_{i}^{t}=\nabla_{x}f_{i}(x_{i}^{t},g(\boldsymbol{x}^{t}))-\nabla g_{i}(x_{i}^{t})\nabla_{y}\bar{f}(\boldsymbol{x}^{t},g(\boldsymbol{x}^{t})).$
According to the definitions of $u_{i}^{t}$ and $\breve{u}_{i}^{t}$, we have
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},$
$\textstyle\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]+\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$ (61)
$\textstyle\leq\frac{\kappa}{2}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{b_{1}}{\kappa(t+1)^{\beta_{1}}}+\frac{b_{2}}{(t+1)^{\beta_{2}}},$
@@
$\vspace{-0.2em}d_{0}$ $\textstyle=\frac{4\bar{L}_{f}^{2}\mathbb{E}[\|\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0}\|^{2}]}{(\lambda_{y}^{0})^{2}(\ln((1-|\rho_{2}|)^{2})e)^{2}};~{}d_{1}\!\!=\!n\sigma_{x}^{2}\!+\!\frac{(\lambda_{x}^{0})^{2}(c_{u1}(C_{z}+r\sigma_{z}^{2})+c_{u2})}{m(\lambda_{z}^{0})^{2}};$
@@
$d_{4}$ $\textstyle=4L_{g}^{2}(\lambda_{x}^{0})^{2}(\lambda_{y}^{0})^{2}(\lambda_{z}^{0})^{-2}(C_{z}c_{u1}+mrc_{u1}\sigma_{z}^{2})$
$\quad+2L_{g}^{2}(\lambda_{y}^{0})^{2}(2C_{x}+n\sigma_{x}^{2}+2mc_{u2}(\lambda_{x}^{0})^{2})$
@@
$d_{5}$ $\textstyle=\frac{2^{2\min\{2v_{x}+2\varsigma_{y},2v_{y}+1,2v_{y}+2\varsigma_{x}\}+2}d_{4}\bar{L}_{f}^{2}}{(\lambda_{y}^{0})^{2}(\ln(1-|\rho_{2}|)e)^{2}|\rho_{2}|}.\vspace{-0.2em}$ (62)
@@
According to Assumption 2, we have $\mathbb{E}[g^{t}(\boldsymbol{x}^{t})-g(\boldsymbol{x}^{t})]=0$. Moreover, the dynamics (32) and Assumption 4 imply $\mathbb{E}[\frac{1}{m}\sum_{i=1}^{m}\tilde{y}_{i}^{t}-g^{t}(\boldsymbol{x}^{t})]=\mathbb{E}[\frac{1}{\lambda_{y}^{t}}\frac{1}{m}\sum_{i=1}^{m}\chi_{wi}^{t}]=0$. Combining these relations and (63) leads to
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$ (64)
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes\frac{1}{m}\sum_{i=1}^{m}\tilde{y}_{i}^{t})\!-\!\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle],$
where in the derivation we have used the fact that $\nabla h(x,y;\varphi_{i})$ is unbiased from Assumption 2.
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$ (66)
$\textstyle=\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\boldsymbol{\hat{y}}^{t+1}-\boldsymbol{\hat{y}}^{t})\rangle].$

### v3

@@ VII-A Auxiliary lemmas
$\textstyle\leq 4m\sigma_{f}^{2}(\lambda_{z}^{t})^{2}t^{-1}+4\bar{L}_{f}^{2}(\lambda_{z}^{t})^{2}\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t-1}\|^{2}]$
@@
Following an argument similar to the derivation of (34), we incorporate the preceding inequality and (43) into the second term on the right hand side of (40) to establish Lemma 7.
@@
Under Assumptions 1-4, if the decaying rates of stepsizes satisfy $1>v_{x}>v_{z}>v_{y}>0$, then we have
$\textstyle\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]$ $\textstyle\leq\frac{\bar{C}_{x}}{(t+1)^{2\varsigma_{x}}},$ (44)
$\textstyle\mathbb{E}[\|\boldsymbol{y}^{t+1}-\boldsymbol{y}^{t}-\boldsymbol{1}_{m}\otimes(\bar{y}^{t+1}-\bar{y}^{t})\|^{2}]$ $\textstyle\leq\frac{\bar{C}_{y}}{(t+1)^{2\varsigma_{y}}},$ (45)
$\textstyle\mathbb{E}[\|\boldsymbol{z}^{t+1}-\boldsymbol{z}^{t}-\boldsymbol{1}_{m}\otimes(\bar{z}^{t+1}-\bar{z}^{t})\|^{2}]$ $\textstyle\leq\frac{\bar{C}_{z}}{(t+1)^{\beta_{z}}},$ (46)
where $\beta_{z}$ is given by $\beta_{z}=\min\{2v_{z}+2\varsigma_{y}-2v_{y},2\varsigma_{z}\}$ and $\bar{C}_{x}$, $\bar{C}_{y}$, and $\bar{C}_{z}$ are given in (50), (51), and (52), respectively.
@@
$\textstyle\sum_{i=1}^{m}\!\!\left(\mathbb{E}[\|\mathbf{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t+1}\|^{2}]\!+\!\frac{|\rho_{2}|}{2}\!\mathbb{E}[\|\mathbf{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]\right)$
@@
$\textstyle\leq\left(1-\frac{|\rho_{2}|}{2}\right)\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]$
$\textstyle+(c_{y1}(\lambda_{y}^{t})^{2}+c_{z1}(\lambda_{z}^{t})^{2})\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t-1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t-1}\|^{2}]$
@@
$\textstyle\sum_{i=1}^{m}\!\!\left(\mathbb{E}[\|\mathbf{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t+1}\|^{2}]\!+\!\frac{|\rho_{2}|}{2}\!\mathbb{E}[\|\mathbf{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]\right)$
@@
$\textstyle\leq\left(1-\frac{|\rho_{2}|}{2}\right)\Big[\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]$
$\textstyle\quad+\frac{|\rho_{2}|}{2}\!\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]$
@@
$\textstyle\sum_{i=1}^{m}\!\!\left(\mathbb{E}[\|\mathbf{x}_{(i)}^{t+1}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t+1}\|^{2}]\!+\!\frac{|\rho_{2}|}{2}\!\mathbb{E}[\|\mathbf{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]\right)$ (48)
@@
where $\beta_{0}=\min\{2v_{z}+2\varsigma_{y}-2v_{y},2\varsigma_{x},2\varsigma_{y},2\varsigma_{z}\}$ and $c_{1}\!=\!b_{0}(\frac{4\beta_{0}}{e\ln(\frac{4}{4-|\rho_{2}|})})^{\beta_{0}}(\frac{a_{0}(2-|\rho_{2}|)}{2b_{0}}\!+\!\frac{4}{|\rho_{2}|})$ with $a_{0}\triangleq\sum_{i=1}^{m}\!(\mathbb{E}[\|\mathbf{x}_{(i)}^{1}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{1}\|^{2}]+\frac{|\rho_{2}|}{2}\mathbb{E}[\|\mathbf{x}_{(i)}^{0}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{0}\|^{2}])+(1-\frac{|\rho_{2}|}{4})(\mathbb{E}[\|\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0}\|^{2}]+\mathbb{E}[\|\boldsymbol{\hat{z}}^{1}-\boldsymbol{\hat{z}}^{0}\|^{2}])$ and $b_{0}=c_{x2}+c_{y3}+c_{z4}+2mr$.
@@
$\textstyle\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t+1}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t+1}\|^{2}]$ (49)
$\textstyle\leq(1-|\rho_{2}|)\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]$
@@
$\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]\leq\frac{\bar{C}_{x}}{(t+1)^{2\varsigma_{x}}},$ (50)
where $\bar{C}_{x}$ is given by $\bar{C}_{x}\!=\!b_{1}(\frac{8\varsigma_{x}}{e\ln(\frac{2}{2-|\rho_{2}|})})^{2\varsigma_{x}}(\frac{a_{1}(1-|\rho_{2}|)}{b_{1}}+\frac{2}{|\rho_{2}|})$ with $a_{1}\!=\!\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{0}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{0}\|^{2}]$, $b_{1}\!=\!\frac{4c_{1}c_{x1}(\lambda_{x}^{0})^{2}}{(4-|\rho_{2}|)(\lambda_{z}^{0})^{2}}+c_{x2}\sigma_{x}^{2}$, $c_{1}$ given in (48), and $c_{x1}$ and $c_{x2}$ given in (24).
@@
$\mathbb{E}[\|\boldsymbol{\hat{y}}^{t+1}-\boldsymbol{\hat{y}}^{t}\|^{2}]\leq\frac{\bar{C}_{y}}{(t+1)^{2\varsigma_{y}}},$ (51)
with $\bar{C}_{y}=b_{2}(\frac{8\varsigma_{y}}{e\ln(\frac{2}{2-|\rho_{2}|})})^{2\varsigma_{y}}(\frac{a_{2}(1-|\rho_{2}|)}{b_{2}}+\frac{2}{|\rho_{2}|})$, $a_{2}\!=\!\textstyle\mathbb{E}[\|\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0}\|^{2}]$, $b_{2}\!=\!\bar{C}_{x}c_{y1}(\lambda_{y}^{0})^{2}+c_{y3}\sigma_{y}^{2}+\frac{4c_{1}c_{y2}(\lambda_{y}^{0})^{2}(\lambda_{x}^{0})^{2}}{(4-|\rho_{2}|)(\lambda_{z}^{0})^{2}}$, $c_{1}$ given in (48), $\bar{C}_{x}$ given in (50), and $c_{y1}$ through $c_{y3}$ given in (31).
@@
$\mathbb{E}[\|\boldsymbol{\hat{z}}^{t+1}-\boldsymbol{\hat{z}}^{t}\|^{2}]\leq\frac{\bar{C}_{z}}{(t+1)^{\min\{2v_{z}+2\varsigma_{y}-2v_{y},2\varsigma_{z}\}}}=\frac{\bar{C}_{z}}{(t+1)^{\beta_{z}}},$ (52)
with $\bar{C}_{z}=b_{3}(\frac{4\beta_{z}}{e\ln(\frac{2}{2-|\rho_{2}|})})^{4\beta_{z}}(\frac{a_{3}(1-|\rho_{2}|)}{b_{3}}+\frac{2}{|\rho_{2}|})$, $a_{3}\!=\!\textstyle\mathbb{E}[\|\boldsymbol{\hat{z}}^{1}-\boldsymbol{\hat{z}}^{0}\|^{2}]$, $b_{3}\!=\!\bar{C}_{x}c_{z1}(\lambda_{z}^{0})^{2}\!+\!\frac{2\bar{C}_{y}c_{z2}(\lambda_{z}^{0})^{2}}{(\lambda_{y}^{0})^{2}}\!+\!\frac{4c_{1}c_{z3}(\lambda_{x}^{0})^{2}}{(4-|\rho_{2}|)}+\frac{c_{z4}(\lambda_{z}^{0})^{2}\sigma_{y}^{2}}{(\lambda_{y}^{0})^{2}}+2mr\sigma_{z}^{2}$, $c_{1}$ given in (48), $\bar{C}_{x}$ given in (50), $\bar{C}_{y}$ given in (51), and $c_{z1}$ through $c_{z3}$ given in (39). ∎
@@
$\textstyle u_{i}^{t}=\nabla_{x}f_{i}^{t}(x_{i}^{t},\tilde{y}_{i}^{t})+\nabla g_{i}^{t}(x_{i}^{t})\tilde{z}_{i}^{t};$ (53)
$\textstyle\breve{u}_{i}^{t}=\nabla_{x}f_{i}(x_{i}^{t},g(\boldsymbol{x}^{t}))+\nabla g_{i}(x_{i}^{t})\nabla_{y}\bar{f}(\boldsymbol{x}^{t},g(\boldsymbol{x}^{t})).$
By defining $\mathbf{\bar{x}}^{t}=\mathrm{col}(\bar{\mathrm{x}}_{(1)}^{t},\cdots,\bar{\mathrm{x}}_{(m)}^{t})$, we have
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},$
$\textstyle\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]+\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$ (55)
$\textstyle\leq\frac{\kappa}{2}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{b_{1}}{\kappa(t+1)^{\beta_{1}}}+\frac{b_{2}}{(t+1)^{\beta_{2}}},$
@@
$d_{0}$ $\textstyle=\frac{4\bar{L}_{f}^{2}\mathbb{E}[\|\boldsymbol{\hat{y}}^{1}-\boldsymbol{\hat{y}}^{0}\|^{2}]}{(\lambda_{y}^{0})^{2}(\ln((1-|\rho_{2}|)^{2})e)^{2}};~d_{1}\!\!=\!n\sigma_{x}^{2}\!+\!\frac{(\lambda_{x}^{0})^{2}(c_{u1}(\bar{C}_{z}+r\sigma_{z}^{2})+c_{u2})}{m(\lambda_{z}^{0})^{2}};$
@@
$d_{4}$ $\textstyle=4L_{g}^{2}(\lambda_{x}^{0})^{2}(\lambda_{y}^{0})^{2}(\lambda_{z}^{0})^{-2}(\bar{C}_{z}c_{u1}+mrc_{u1}\sigma_{z}^{2})$
$\quad+2L_{g}^{2}(\lambda_{y}^{0})^{2}(2\bar{C}_{x}+n\sigma_{x}^{2}+2mc_{u2}(\lambda_{x}^{0})^{2})$
@@
$d_{5}$ $\textstyle=\frac{2^{2\min\{2v_{x}+2\varsigma_{y},2v_{y}+1,2v_{y}+2\varsigma_{x}\}+2}d_{4}\bar{L}_{f}^{2}}{(\lambda_{y}^{0})^{2}(\ln(1-|\rho_{2}|)e)^{2}|\rho_{2}|}.$ (56)
@@
According to Assumption 2, we have $\mathbb{E}[g^{t}(\boldsymbol{x}^{t})-g(\boldsymbol{x}^{t})]=0$. Moreover, Assumption 4 imply $\mathbb{E}[\frac{1}{m}\sum_{i=1}^{m}\tilde{y}_{i}^{t}-g^{t}(\boldsymbol{x}^{t})]=0$. Combining these relations and (57) leads to
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$ (58)
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes\frac{1}{m}\sum_{i=1}^{m}\tilde{y}_{i}^{t})\!-\!\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle],$
where in the derivation we have used Assumption 2.
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla_{x}\boldsymbol{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla_{x}\boldsymbol{f}^{t}(\boldsymbol{x}^{t},\boldsymbol{\tilde{y}}^{t})\rangle]$ (60)
$\textstyle=\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{y}^{t}}M_{\boldsymbol{\tau}^{t}}(\boldsymbol{\hat{y}}^{t+1}-\boldsymbol{\hat{y}}^{t})\rangle].$